OpenStudy (supercalifragalisticexspeaalli):

Solve 5x2 = −30x − 65.

OpenStudy (supercalifragalisticexspeaalli):

@Anonymouse2015 @annah.bee @rockhauler98

OpenStudy (supercalifragalisticexspeaalli):

plzz help will fan and medal

OpenStudy (supercalifragalisticexspeaalli):

@XhawaiiangirlX

OpenStudy (rockhauler98):

divide every term by 5 and rewrite the equation x2+6x+13=0 now use the general quadratic formula or completing the square

OpenStudy (rockhauler98):

your problem has a complex roots

OpenStudy (rockhauler98):

5x2 = -30x - 65 5x^2+30x-65=0 divide it by 5 So, x^2+6x-13=0

OpenStudy (supercalifragalisticexspeaalli):

e 5x2 = −30x − 65. x = −3 ± 2i x = −3 ± 4i x = −6 ± 2i x = −6 ± 4i

OpenStudy (rockhauler98):